# SpatioTemporal Difference Network for Video Depth Super-Resolution

### v1

@@ Introduction
Depth data constitutes a fundamental component in various fields, including 3D reconstruction (Im et al. 2018; Yan et al. 2025), virtual reality (Yan et al. 2022b; Zhou et al. 2023), and augmented reality (Song et al. 2020; Yin et al. 2023). Recently, numerous depth super-resolution (DSR) methods (Guo et al. 2018; Wang et al. 2023a, b) have been proposed to reconstruct high-resolution (HR) depth from low-resolution (LR) inputs, achieving remarkable performance. Furthermore, (Sun et al. 2023) introduce a video depth super-resolution (VDSR) framework that effectively aggregates multi-frame RGB-D features, demonstrating substantial improvements over single-frame approaches.
@@
Figure 2: Visualization of (a) RGB and GT depth at frame $t$, (b) absolute difference representations (top) and corresponding histogram distribution (bottom) between GT depth and bicubic-upsampled LR depth (Bic.). (c) shows the error analysis between consecutive frames ($t$ and $t-1$ bicubic-upsampled LR depth), while (d) presents cross frame results between frames $t$ and $t-2$.
@@
Owing to these innovative designs, our method successfully restores accurate HR depth videos. As shown in Fig. 1, STDNet outperforms five state-of-the-art approaches by 32.6% (×4$\times$ 4× 4), 28.8% (×8$\times$ 8× 8), and 27.6% (×16$\times$ 16× 16) in average.
@@
Recently, single-frame DSR methods (Ye et al. 2020; De Lutio et al. 2022; Chen et al. 2024) have made remarkable progress. Existing approaches can be broadly categorized into filtering-based methods (Metzger, Daudt, and Schindler 2023; Zhong et al. 2023; Wang et al. 2024), multi-modal fusion-based methods (Zhong et al. 2021; Wang et al. 2022; Zhao et al. 2023), multi-task collaborative methods (Sun et al. 2021; Yan et al. 2022a), and structure-oriented methods (Wang, Yan, and Yang 2024; Zheng, Han, and Shen 2025). For example, (Kim, Ponce, and Ham 2021) integrate deformable networks with joint image filtering to adaptively transfer RGB information to depth features. To effectively aggregate RGB-D information, (Zhao et al. 2022) introduce a discrete cosine network to disentangle the shared and private information present in RGB and depth features. Additionally, (Sun et al. 2021) and (Tang et al. 2021) employ an auxiliary depth estimation network to effectively fuse RGB and depth features. More recently, several methods have focused on reconstructing high-frequency information. For instance, (Yuan et al. 2023) develop a recurrent structure attention to perform frequency decomposition and edge restoration. Unlike these single-frame approaches, our method focuses more on addressing the spatiotemporal long-tailed distribution issues inherent in VDSR, enabling robust reconstruction of temporally consistent HR depth videos.
@@
Single-modal video RGB super-resolution (VSR) aims to restore HR RGB videos from the corresponding LR inputs. Existing VSR methods (Isobe et al. 2022; Shi et al. 2022) can be categorized into sliding window-based approaches and recurrent-based approaches. Specifically, sliding window-based methods (Wang et al. 2019; Li et al. 2020; Cao et al. 2021) employ a temporal sliding window strategy to align neighboring frames with the current frame within the window, achieving impressive performance. However, such methods are typically limited by their fixed window size and can only integrate information from a restricted number of nearby frames. To overcome this constraint, recent work has introduced bidirectional recurrent architectures (Chan et al. 2021, 2022; Zhou et al. 2024) that can aggregate information across the entire video sequence for more comprehensive feature enhancement. In contrast to single-modal VSR, VDSR presents unique challenges in establishing multi-frame and multi-modal correspondences between RGB and depth videos. These fundamental discrepancy hinder the effectiveness of existing advanced VSR methods in reconstructing depth videos.
@@
Although single-frame depth enhancement methods have demonstrated success in recovering high-quality depth from degraded inputs, they often exhibit suboptimal performance and temporal inconsistency when applied to depth video sequences. To address these limitations, recent advances in video depth enhancement have investigated temporal fusion strategies that integrate information from both current and adjacent frames. For example, (Sun et al. 2023) propose the first dToF-based VDSR method, which effectively harnesses rich information from multi-frame videos sequences to mitigate spatial ambiguity in LR depth videos. In addition, (Dong et al. 2024) develop a multi-frame depth denoising network that models the intra-scene geometric correlations and inter-frame noise distribution correlations, effectively suppressing multi-path interference and shot noise. More recently, (Zhu et al. 2025) introduce a video depth completion framework that integrates multi-frame features through an adaptive frequency selection fusion module. Different from them, we focus on leveraging the long-tailed distribution characteristics of depth videos to enhance both non-smooth regions and temporal variation zones.
@@
Figure 4: Details of (a) spatial difference, and (b) histogram comparison between our STDNet and DVSR (Sun et al. 2023).
@@
where $f_{bd}$ and $f_{bu}$ are bilinear downsampling and upsampling operations (×2$\times$ 2× 2), respectively. Subsequently, the RGB feature $\boldsymbol{F}_{r}$, depth feature $\boldsymbol{F}_{d}$, and $\boldsymbol{\sigma}$ are fed into the proposed spatial difference for intra-frame RGB-D aggregation.
@@
where $\mathcal{E}_{w}$ consists of a $\times 3$ 3 convolutional layer, maximum function, mean function, and sigmoid function.
@@
Given $\boldsymbol{F}_{r}$, $\boldsymbol{w}_{t}$, $\boldsymbol{F}_{sd}$, $\boldsymbol{\varphi}$ and $\widehat{\boldsymbol{\varphi}}$, we then follow a common bidirectional iterative scheme (Chan et al. 2022; Sun et al. 2023) to execute the proposed temporal difference strategy, fully aggregating information from adjacent RGB and depth frames. Please see our appendix for more iterative details.
@@
Figure 5: Details of (a) temporal difference, and (b) temporal consistency visualization for $x$$t$ slices (along dashed line). Diff.: Difference. Orange rectangular boxes are the deformable convolutional layers (Zhu et al. 2019).
Methods ×4$\times$ 4× 4 ×8$\times$ 8× 8 ×16$\times$ 16× 16 Venue
@@
Table 1: Quantitative comparisons on the TarTanAir. The parameters are computed at the ×8$\times$ 8× 8 scaling factor.
Temporal Difference. Our temporal difference is delineated in Fig. 5(a), consists of adjacent frame fusion and cross frame fusion. Taking forward propagation as an example, we use RGB and depth features from adjacent frames ($t-1$ and $t-2$) to enhance temporal variation regions in the $t$-th frame depth, guided by temporal difference representations $\boldsymbol{\varphi}_{t-1}$ and $\widehat{\boldsymbol{\varphi}}_{t-2}$. Conversely, the backward propagation employs subsequent RGB-D frames ($t+1$ and $t+2$) for complementary refinement of $t$-th frame depth.
In the adjacent frame fusion stage, we utilize a temporal difference encoder $\mathcal{E}_{\varphi}$ (composed of convolutional layers and sigmoid function) to project $\boldsymbol{\varphi}_{t-1}$ into offset $\delta_{t-1}$ and modulation scalar $m_{t-1}$. Deformable convolution $\mathcal{D}$ (Zhu et al. 2019) is then introduced to dynamically sample temporal variation information matched with the temporal difference representations. Additionally, adaptive weight $\boldsymbol{w}_{t}$ derived from the spatial difference branch is employed to mitigate cross-modal discrepancies between adjacent RGB frames and the current depth frames. The enhanced depth feature $\boldsymbol{F}_{f}^{t-1,t}$ for frame $t$ can be expressed as:
@@
where intermediate feature $\boldsymbol{F}_{f,dc}^{t-1}=\mathcal{D}(\boldsymbol{F}_{f}^{t-1},\delta_{t-1},m_{t-1})$, $\boldsymbol{F}_{r,dc}^{t-1}=\mathcal{D}(\boldsymbol{F}_{r}^{t-1},\delta_{t-1},m_{t-1})$. $\boldsymbol{F}_{f}^{t}$ and $\boldsymbol{F}_{f}^{t-1}$ are the depth features of frame $t$ and frame $t-1$ during forward iteration.
@@
Figure 6: Visual results of consecutive frames on the TarTanAir at ×16$\times$ 16× 16 upscaling.
Figure 7: Visual results on the DyDToF and DynamicReplica at ×8$\times$ 8× 8 upscaling.
@@
Given the GT depth video $\boldsymbol{D}_{GT}$ and predicted HR depth video $\boldsymbol{D}_{HR}$, we follow prior work (Sun et al. 2023) by applying Charbonnier regularization (Charbonnier et al. 1994) as reconstruction loss $\mathcal{L}_{rec}$ to constrain our STDNet:
@@
For the spatial difference term, we introduce an uncertainty constraint (Ning et al. 2021) to facilitate the learning of difference representations in non-smooth regions:
$\mathcal{L}_{sd}={\textstyle\sum_{q\in\mathbb{Q}}}(\boldsymbol{\sigma}^{q}-min(\boldsymbol{\sigma}^{q}))||\boldsymbol{D}_{GT}^{q}-\boldsymbol{D}_{HR}^{q}||_{1},$ (11)
@@
$\begin{split}&\mathcal{L}_{td}=\underbrace{{\textstyle\sum_{q\in\mathbb{Q}}}||\mathcal{R}_{df}(\boldsymbol{\varphi}^{q})-\Phi(\boldsymbol{D}_{GT}^{q})||_{1}}_{\text{adjacent frame}}\\ &+\underbrace{{\textstyle\sum_{q\in\mathbb{Q}}}||\mathcal{R}_{df}(\widehat{\boldsymbol{\varphi}}^{q})-\Phi(\boldsymbol{D}_{GT}^{q})||_{1}}_{\text{cross frame}},\end{split}$ (12)
@@
Datasets. Following previous methods (Sun et al. 2023; Zhu et al. 2025), we evaluate STDNet on TarTanAir (Wang et al. 2020), DyDToF (Sun et al. 2023), and DynamicReplica (Karaev et al. 2023) datasets. Since dataset pre-processing details from prior approaches are unavailable, we redefine the training and test sets. Specifically, we utilize the hard scenes from TarTanAir for training, consisting of 6,1646,1646 , 164 RGB-D frames in the train set and 1,2281,2281 , 228 frames in the test set. Then, the pre-trained model on TarTanAir is evaluated on DyDToF (576576576 frames) and DynamicReplica (500500500 frames) without any fine-tuning. Besides, DynamicReplica is center-cropped to match the size of TarTanAir (640×480640$\times$ 480640 × 480). All LR depth are generated from GT and RGB using the same synthesis pipeline as DVSR (Sun et al. 2023). We retrain the existing methods from scratch using their released code.
Implementation Details. Consistent with prior works (Sun et al. 2023; Zhu et al. 2025), we adopt root mean square error (RMSE), mean absolute error (MAE), and temporal end-point error (TEPE) as evaluation metrics (all measured in centimeters). During training, we randomly crop RGB and HR depth frames to $\times 256$ 256. Besides, we employ the Adam optimizer (Kingma 2014) with an initial learning rate of $1\times 10^{-4}$ to train our STDNet. Our method is implemented using two NVIDIA RTX 4090 GPUs. The hyper-parameters are set as $\alpha_{1}$=$\alpha_{2}$=0.5 and $\beta$=0.01.
@@
We compare our STDNet with existing state-of-the-art approaches, including DJFR (Li et al. 2019), CUNet (Deng and Dragotti 2020), DKN (Kim, Ponce, and Ham 2021), FDSR (He et al. 2021), SUFT (Shi, Ye, and Du 2022), SGNet (Wang, Yan, and Yang 2024), C2PD (Kang et al. 2025), DORNet (Wang et al. 2025), and DVSR (Sun et al. 2023).
Methods ×4$\times$ 4× 4 ×8$\times$ 8× 8 ×16$\times$ 16× 16
@@
[…]
Methods ×4$\times$ 4× 4 ×8$\times$ 8× 8 ×16$\times$ 16× 16
@@
Visual Comparison. Figs. 6 and  7 provide visual comparisons, clearly indicating that our method achieves more accurate depth recovery. For example, compared to previous approaches, the structure and shape of chair in Fig. 6 predicted by STDNet align more closely with the GT depth while exhibiting superior temporal stability. Additionally, Fig. 7 shows that our method results in more precise reconstruction of dynamic objects (e.g., dog and cat) than others.
[…]
Figure 8: Complexity on the TarTanAir tested by a 4090 GPU. A larger circle area indicates larger time.
@@
[…]
@@
Figure 9: Ablation study of spatial difference (SD) and temporal difference (TD) on the TarTanAir and DyDToF (×4$\times$ 4× 4).
Figure 10: Ablation study of STDNet with (a) different numbers of adjacent frames and different loss functions (×16$\times$ 16× 16).
Different Numbers of Adjacent Frames. Fig. 10(a) reveals the impact of different numbers of neighboring frames in TD. Compared to using only one RGB-D frames, incorporating additional neighboring frames significantly reduces the RMSE. However, the performance gains diminish when the number exceeds 222. Consequently, to balance computational cost and performance, our STDNet selects 222 frames (a adjacent frame and a cross frame) as the default setting.

### v2

@@ Introduction
Depth data constitutes a fundamental component in various fields, including 3D reconstruction (im2018accurate; lian2025cross; yan2025completion), virtual reality (yan2022rignet; lian2023multi; zhou2023memory; yan2024learnable), and augmented reality (song2020channel; yan2025event; yin2023metric3d). Recently, numerous depth super-resolution (DSR) methods (guo2018hierarchical; wang2023rgb; wang2023learning) have been proposed to reconstruct high-resolution (HR) depth from low-resolution (LR) inputs, achieving remarkable performance. Furthermore, (sun2023consistent) introduce a video depth super-resolution (VDSR) framework that effectively aggregates multi-frame RGB-D features, demonstrating substantial improvements over single-frame approaches.
@@
Figure 2: Visualization of (a) RGB and GT depth at frame $t$, (b) absolute difference representations (top) and corresponding histogram distribution (bottom) between GT depth and bicubic-upsampled LR depth (Bic.). (c) shows the error analysis between consecutive frames ($t$ and t1t-1 bicubic-upsampled LR depth), while (d) presents cross frame results between frames $t$ and t2t-2.
@@
Owing to these innovative designs, our method successfully restores accurate HR depth videos. As shown in Fig. 1, STDNet outperforms five state-of-the-art approaches by 32.6% ($\times 4$), 28.8% ($\times 8$), and 27.6% ($\times 16$) in average.
@@
Recently, single-frame DSR methods (ye2020pmbanet; de2022learning; chen2024intrinsic) have made remarkable progress. Existing approaches can be broadly categorized into filtering-based methods (metzger2023guided; zhong2023deep; wang2024scene), multi-modal fusion-based methods (zhong2021high; wang2022depth; zhao2023spherical), multi-task collaborative methods (sun2021learning; yan2022learning), and structure-oriented methods (yan2025ducos; bi2025dual; zheng2025decoupling; bi2025structure). For example, (kim2021deformable) integrate deformable networks with joint image filtering to adaptively transfer RGB information to depth features. To effectively aggregate RGB-D, (zhao2022discrete) introduce a discrete cosine network to disentangle the shared and private information present in RGB and depth features. Additionally, (sun2021learning) and (tang2021bridgenet) employ an auxiliary depth estimation network to effectively fuse RGB and depth features. More recently, several methods have focused on reconstructing high-frequency information. For instance, (yuan2023recurrent) develop a recurrent structure attention to perform frequency decomposition and edge restoration. Unlike these single-frame approaches, our method focuses more on addressing the spatiotemporal long-tailed distribution issues inherent in VDSR, enabling robust reconstruction of temporally consistent HR depth videos.
@@
Single-modal video RGB super-resolution (VSR) aims to restore HR RGB videos from the corresponding LR inputs. Existing methods (isobe2022look; gao2022lightweight; shi2022rethinking) can be categorized into sliding window-based and recurrent-based approaches. Specifically, sliding window-based methods (wang2019edvr; li2020mucan; cao2021video) employ a temporal sliding window to align neighboring frames with the current frame within the window, achieving impressive performance. However, such methods are typically limited by their fixed window size and can only integrate information from a restricted number of nearby frames. To overcome this constraint, recent work has introduced bidirectional recurrent architectures (chan2021basicvsr; hu2025exploiting; zhou2024video) that can aggregate information across the entire video sequence for feature enhancement. In contrast to VSR, VDSR presents unique challenges in establishing multi-frame and multi-modal correspondences between RGB and depth videos. These fundamental discrepancy hinder the effectiveness of existing advanced VSR methods in reconstructing depth videos.
@@
Although single-frame depth enhancement methods have demonstrated success in recovering high-quality depth from degraded inputs, they often exhibit suboptimal performance and temporal inconsistency when applied to depth video sequences. To address these limitations, recent advances in video depth enhancement have investigated temporal fusion strategies that integrate information from both current and adjacent frames. For example, (sun2023consistent) propose the first dToF-based VDSR method, which effectively harnesses rich information from multi-frame videos sequences to mitigate spatial ambiguity in LR depth videos. In addition, (dong2024exploiting) develop a multi-frame depth denoising network that models the intra-scene geometric correlations and inter-frame noise distribution correlations, effectively suppressing multi-path interference and shot noise. More recently, (zhu2025svdc) introduce a video depth completion framework that integrates multi-frame features through an adaptive frequency selection fusion module. Different from them, we focus on leveraging the long-tailed distribution characteristics of depth videos to enhance both non-smooth regions and temporal variation zones.
@@
Figure 4: Details of (a) spatial difference, and (b) histogram comparison between our STDNet and DVSR (sun2023consistent).
@@
where $f_{bd}$ and $f_{bu}$ are bilinear downsampling and upsampling operations ($\times 2$), respectively. Subsequently, the RGB feature $\boldsymbol{F}_{r}$, depth feature $\boldsymbol{F}_{d}$, and $\boldsymbol{\sigma}$ are fed into the proposed spatial difference for intra-frame RGB-D aggregation.
@@
where $\mathcal{E}_{w}$ consists of a $3\times 3$ convolutional layer, maximum function, mean function, and sigmoid function.
@@
Given $\boldsymbol{F}_{r}$, $\boldsymbol{w}_{t}$, $\boldsymbol{F}_{sd}$, $\boldsymbol{\varphi}$ and $\widehat{\boldsymbol{\varphi}}$, we then follow a common bidirectional iterative scheme (chan2022basicvsr++; sun2023consistent) to execute the proposed temporal difference strategy, fully aggregating information from adjacent RGB and depth frames. Please see our appendix for more iterative details.
@@
Figure 5: Details of (a) temporal difference, and (b) temporal consistency visualization for xxtt slices (along dashed line). Diff.: Difference. Orange rectangular boxes are the deformable convolutional layers (zhu2019deformable).
Methods $\times 4$ $\times 8$ $\times 16$ Venue
@@
Table 1: Quantitative comparisons with existing state-of-the-art methods on the TarTanAir dataset.
Temporal Difference. Our temporal difference is delineated in Fig. 5(a), consists of adjacent frame fusion and cross frame fusion. Taking forward propagation as an example, we use RGB and depth features from adjacent frames (t1t-1 and t2t-2) to enhance temporal variation regions in the $t$-th frame depth, guided by temporal difference representations $\boldsymbol{\varphi}_{t-1}$ and $\widehat{\boldsymbol{\varphi}}_{t-2}$. Conversely, the backward propagation employs subsequent RGB-D frames ($t+1$ and $t+2$) for complementary refinement of $t$-th frame depth.
In the adjacent frame fusion stage, we utilize a temporal difference encoder $\mathcal{E}_{\varphi}$ (composed of convolutional layers and sigmoid function) to project $\boldsymbol{\varphi}_{t-1}$ into offset $\delta_{t-1}$ and modulation scalar $m_{t-1}$. Deformable convolution $\mathcal{D}$ (zhu2019deformable) is then introduced to dynamically sample temporal variation information matched with the temporal difference representations. Additionally, adaptive weight $\boldsymbol{w}_{t}$ derived from the spatial difference branch is employed to mitigate cross-modal discrepancies between adjacent RGB frames and the current depth frames. The enhanced depth feature $\boldsymbol{F}_{f}^{t-1,t}$ for frame $t$ can be expressed as:
@@
where intermediate feature $\boldsymbol{F}_{f,dc}^{t-1}=\mathcal{D}(\boldsymbol{F}_{f}^{t-1},\delta_{t-1},m_{t-1})$, $\boldsymbol{F}_{r,dc}^{t-1}=\mathcal{D}(\boldsymbol{F}_{r}^{t-1},\delta_{t-1},m_{t-1})$. $\boldsymbol{F}_{f}^{t}$ and $\boldsymbol{F}_{f}^{t-1}$ are the depth features of frame $t$ and frame t1t-1 during forward iteration.
@@
Figure 6: Visual results of consecutive frames on the TarTanAir at $\times 16$ upscaling.
Figure 7: Visual results on the DyDToF at $\times 8$ upscaling.
@@
Given the GT depth video $\boldsymbol{D}_{GT}$ and predicted HR depth video $\boldsymbol{D}_{HR}$, we follow prior work (sun2023consistent) by applying Charbonnier regularization (charbonnier1994two) as reconstruction loss $\mathcal{L}_{rec}$ to constrain our STDNet:
@@
For the spatial difference term, we introduce an uncertainty constraint (ning2021uncertainty) to facilitate the learning of difference representations in non-smooth regions:
[…]
@@
[…]
@@
Datasets. Following previous methods (sun2023consistent; zhu2025svdc), we evaluate STDNet on TarTanAir (wang2020TarTanAir), DyDToF (sun2023consistent), and DynamicReplica (karaev2023dynamicstereo) datasets. Since dataset pre-processing details from prior approaches are unavailable, we redefine the training and test sets. Specifically, we utilize the hard scenes from TarTanAir for training, consisting of $6,164$ RGB-D frames in the train set and $1,228$ frames in the test set. Then, the pre-trained model on TarTanAir is evaluated on DyDToF ($576$ frames) and DynamicReplica ($500$ frames) without any fine-tuning. Besides, DynamicReplica is center-cropped to match the size of TarTanAir ($640\times 480$). All LR depth are generated from GT and RGB using the same synthesis pipeline as DVSR (sun2023consistent). We retrain the existing methods from scratch using their released code.
Implementation Details. Consistent with prior works (sun2023consistent; zhu2025svdc), we adopt root mean square error (RMSE), mean absolute error (MAE), and temporal end-point error (TEPE) as evaluation metrics (all measured in centimeters). During training, we randomly crop RGB and HR depth frames to $256\times 256$. Besides, we employ the Adam optimizer (kingma2014adam) with an initial learning rate of $1\times 10^{-4}$ to train our STDNet. Our method is implemented using two NVIDIA RTX 4090 GPUs. The hyper-parameters are set as $\alpha_{1}$$\alpha_{2}$=0.5 and $\beta$=0.01.
@@
We compare our STDNet with existing state-of-the-art approaches, including DJFR (li2019joint), CUNet (deng2020deep), DKN (kim2021deformable), FDSR (he2021towards), SUFT (shi2022symmetric), SGNet (wang2024sgnet), C2PD (kang2025c2pd), DORNet (wang2025dornet), and DVSR (sun2023consistent).
Methods $\times 4$ $\times 8$ $\times 16$
@@
Quantitative Comparison. Tabs. 1-3 list quantitative comparisons across multiple datasets, demonstrating that our STDNet outperforms existing state-of-the-art approaches at $\times 4$, $\times 8$, and $\times 16$ scaling factors. Specifically, Tab. 1 shows that our method is superior to both the existing multi-frame DVSR and advanced single-frame approaches on TarTanAir. For example, compared to the second-best method, our STDNet reduces RMSE by15.24cm15.24cm, MAE by $2.38cm$, and TEPE by $2.16cm$ on $\times 16$ VDSR. Additionally, Tabs. 2 and 3 further validate the generalization capability of our method on DyDToF and DynamicReplica. We observe that STDNet achieves outstanding performance, surpassing the suboptimal VDSR ($\times 16$) by $4.31cm$ in RMSE on DyDToF and by $0.15cm$ in RMSE on DynamicReplica.
Methods $\times 4$ $\times 8$ $\times 16$
@@
Visual Comparison. Figs. 6 and  7 provide visual comparisons, clearly indicating that our method achieves more accurate depth recovery. For example, compared to previous approaches, the structure and shape of chair in Fig. 6 predicted by STDNet align more closely with the GT depth while exhibiting superior temporal stability. Additionally, Fig. 7 shows that our method results in more precise reconstruction of dynamic objects (e.g., dog) than others.
Model Complexity Analysis. Fig. 8 demonstrates that our STDNet maintains a comparable balance between parameters, performance, and inference time. Compared to single-frame approaches (DKN, SUFT, SGNet, C2PD, and DORNet), although our method exhibits higher time cost, it achieves a significant average reduction of $9.23M$ parameters and $35.82cm$ RMSE. Furthermore, STDNet outperforms the multi-frame DVSR with a $47.35ms$ improvement in inference speed and a $4.93cm$ gain in performance, while introducing only a modest parameter increase of $4.4M$.
Figure 8: Complexity on the TarTanAir ($\times 8$) tested by a 4090 GPU. A larger circle area indicates larger time.
@@
Spatial Difference and Temporal Difference. Fig. 9 illustrates an ablation study of spatial difference (SD) and temporal difference (TD). For the baseline, we replace all SD and TD in STDNet with concatenation operations, while keeping other architecture unchanged. As shown in Fig. 9(a), both SD and TD contribute to performance improvements over the baseline. For example, SD and TD reduce RMSE by $3.56cm$ and $14.02cm$ respectively on the DyDToF. When deployed together, STDNet achieves the best performance, surpassing the baseline by $17.94cm$ on the DyDToF.
@@
Figure 9: Ablation study of spatial difference (SD) and temporal difference (TD) on the TarTanAir and DyDToF ($\times 4$).
Figure 10: Ablation study of STDNet with (a) different numbers of adjacent frames and different loss functions ($\times 16$).
Different Numbers of Adjacent Frames. Fig. 10(a) reveals the impact of different numbers of neighboring frames in TD. Compared to using only one RGB-D frames, incorporating additional neighboring frames significantly reduces the RMSE. However, the performance gains diminish when the number exceeds $2$. Consequently, to balance computational cost and performance, our STDNet selects $2$ frames (a adjacent frame and a cross frame) as the default setting.